OpenStudy (anonymous):

Anthropologists can estimate the height of a woman by measuring the length of her radius bone (from the wrist to the elbow). The length of the radius bone b is given by b=0.26h-18.85 where h is the height (in centimeters) of the woman. How do you solve the equation?

OpenStudy (anonymous):

If you are given any information h or b in the problem, you can just plug those into the equation and solve for the other variable. But to find the general equation for height of a woman (ie find h in terms of b), then you can just solve the equation for h (isolate h).

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

If they tell you what h or b is, plug in that information and solve the equation. If they ask to solve in terms of h, then make the equation into another equal equation that starts with "h=". From the information given, I don't see how you can solve this except for solving it in terms of h.

OpenStudy (anonymous):

But i dont get how you would solve for h..

OpenStudy (anonymous):

Do you know how to solve 3x+1=10 solved for x lead to 3x=9 which leads to x=3. You do the opposite of what things other than the variable you're solving for are doing to the variable. So a 3 is being multiplied by the x and the x is being subtracted by 1. You work in revers order of operations (PEMDAS). So if we have 2 variables like 3x+y=10, then to solve for x we would just treat the y as something being done to the x. So it would look like 3x+y=10 3x=10-y (because the x was being added by y, and we did the opposite to isolate x on one side of the equation) \[x=\frac{ 10-y }{ 3 }\] (because the x was being multiplied by the 3, we divide to cancel it out on that side) And this answer is 3x+y=10 solved in terms of x.